find the equation of the line through (3,5)(8,15) write the answer in standard form using only integers
Answer by edjones(8007)

(3,5)(8,15)
m=(15-5)/(8-3)
=10/5
=2 slope
y-5=2(x-3)
y-5=2x-6
y=2x-1
-2x+y=-1 Standard form.
.
Ed